3. You place 50.0 mL of 0.500 M NaOH in a coffee-cup calorimeter at 35.00à ƒ ƒ ¡ µ ’C and add 25.00 mL of
0.500 M HCl, also at 35.00à ƒ ƒ ¡ µ ’C. After stirring, the final temperature is 37.21à ƒ ƒ ¡ µ ’C. [Assume the total
volume is the sum of the individual volumes and that the final solution has the same density (1.00
g/mL) and specific heat capacity as water (4.184 J/g.K).
(a). Calculate the qsoln (in J and kJ)
(b) Calculate the change in enthalpy, à ƒ ƒ ¢ ˆ †H of the reaction (in kJ/mol of water formed)
1
Expert's answer
2021-10-17T11:24:54-0400

Molar mass of NaOH = 39.997 g/mol

39.997×0.5= 19.9985

50/19.9985= 2.5M

Molar mass of HCl= 36.458 g/mol

36.458×0.5= 18.229

25/18.229= 1.37M

37.21-35= 2.21

2.21+273= 275.21k

= 4.184×275.21×3.87

= 4456.22J